OpenStudy (anonymous):

Solve the equation: (x + 1)2 = 9. Enter the largest zero value.

OpenStudy (nenadmatematika):

(x+1)=3 or (x+1)=-3 so x=2 or x=-4

OpenStudy (earthcitizen):

x=7/2

OpenStudy (nenadmatematika):

? (7/2+1)^2=81/4

OpenStudy (earthcitizen):

did u mean \[(x+1)^{2}=9\]

OpenStudy (anonymous):

it would be 7/2 if i was just solving the problem, but i have to find the zeros? so its a different way to solve it.

OpenStudy (phi):

@Earth air means (x+1)^2 It is a bad habit people use here.

OpenStudy (earthcitizen):

yh, i dig. thnx, thought it was (x+1)2